OpenStudy (witz):

What is the vertex form of the equation? y = x^2 – 8x + 12

OpenStudy (loviy):

Try to use the form y= (x - h)^2 + k where (h, k) is the vertex.

OpenStudy (loviy):

The goal here is to use Completing the square.

OpenStudy (loviy):

For example, if we have y = x^2 - 4x + 1. We will take half of the coefficient of x (of half of -4), and then square it.

OpenStudy (loviy):

So, half of -4 is -2 and squaring it is 4.. We need to add and subtract that... y = x^2 - 4x + 4 + 1 - 4. Then, y = (x^2 - 4x + 4) + 1 - 4.

OpenStudy (loviy):

If we will factor x^2 - 4x + 4 it would be (x - 2)(x - 2) or (x - 2)^2. So, it would be y = (x - 2)^2 - 3. Pleasure to help you!

OpenStudy (lyss):

yep ignore me.

OpenStudy (witz):

thank you, loviy! this explanation's very helpful.

OpenStudy (loviy):

You're welcome! God bless you! :)